OpenStudy (anonymous):

Why is the exponent 0 of any number equal to 1?

OpenStudy (anonymous):

by definition

OpenStudy (ash2326):

let's say we have a^2/a^2= a^(2-2) a^0=a^2/a^2 a^0=1

OpenStudy (anonymous):

it is the definition that will extend the laws of exponents to zero and negative numbers. so for example \[10^3=1000, 10^2=100, 10^2=10\] so it makes sense to say \[10^0=1\] to complete the pattern of dividing by ten

OpenStudy (binary3i):

guys started to plot graph they begain from -infinity to infinity but were stuck at 0 thy abserved the limits and it turned out to be one

OpenStudy (turingtest):

Actually it's not true for 0^0 that form is indeterminate.

OpenStudy (anonymous):

what turingtest said

OpenStudy (anonymous):

although theres is apparently still some "controversy" about \[0^0\]

OpenStudy (turingtest):

True, it's a bit debated.

OpenStudy (anonymous):

Thank you everyone.